# Explaining Sources of Uncertainty in Automated Fact-Checking

### v1

@@ 5.1 Faithfulness
To assess whether the NLEs produced by CLUE are faithful to the model’s uncertainty, we adapt the Correlational Counterfactual Test (CCT) Siegel et al. (2024) and propose an Entropy-CCT metric.
Following Siegel et al. (2024), we begin by inserting a random adjective or noun into the original instance $X$ to obtain a perturbed input $X^{\prime}$ (See App. D for details). Let $u(X)$ denote the model’s uncertainty score defined by Eq. 2, unlike CCT(See details of original CCT in App.E), we measure the impact of the perturbation on the model’s uncertainty with Absolute Entropy Change (AEC):
$\Delta u(X)=\left|u(X)-u(X^{\prime})\right|$ (8)
For each perturbation, we record whether the inserted word appears in the generated NLE, using its presence as a proxy for importance. This yields a binary mention flag $m\in\{0,1\}$, following Siegel et al. (2024); Atanasova et al. (2023).
Let $D_{m}$ denote the set of perturbed examples where the NLE mentions the inserted word and $D_{\lnot m}$ is the complementary set where it does not, we correlate the continuous variable $\Delta u$ with the binary mention flag $m$ via the point-biserial correlation $r_{\text{pb}}$ (Tate, 1954). The Entropy-CCT statistic is:
$\scriptstyle\text{CCT}_{\text{entropy}}=r_{\text{pb}}=\frac{\mathbb{E}_{m}[\Delta u]-\mathbb{E}_{\lnot m}[\Delta u]}{\mathrm{Std}(\Delta u)}\cdot\sqrt{\frac{|D_{m}|\cdot|D_{\lnot m}|}{(|D_{m}|+|D_{\lnot m}|)^{2}}}$ (9)
where $\mathbb{E}_{m}[\Delta u]$ and $\mathbb{E}_{\lnot m}[\Delta u]$ are the mean absolute entropy changes for these two groups, respectively. $\mathrm{Std}(\Delta u)$ is the standard deviation of absolute entropy changes across the full dataset.
Ultimately, this metric quantifies the alignment between changes in model uncertainty and explanatory references to input perturbations, thereby measuring how faithfully the NLEs reflect the model’s uncertainty.
@@ 5.2 Span-Coverage
[…]
@@ 5.3 Span-Extraneous
[…]
@@ 5.4 Label-Explanation Entailment
We evaluate the extent to which the uncertainty explanation agrees with the model’s predicted label by formulating the task as a natural-language inference (NLI) problem. First, we convert the predicted label into a hypothesis using the template “The claim is supported by / refuted by / neutral to the evidence.” The generated explanation serves as the premise. The resulting premise–hypothesis pair is fed to a widely used off-the-shelf language-inference model, DeBERTa-v3666https://huggingface.co/MoritzLaurer/DeBERTa-v3-large-mnli-fever-anli-ling-wanli (He et al., 2023). The Label-Explanation Entailment (LEE) score is the proportion of examples for which the NLI model predicts entailment.
@@ 5.5 Results
HealthVer DRUID
Approach Faith. ($\uparrow$) Span-Cov. ($\uparrow$) Span-Ext. ($\downarrow$) LEE ($\uparrow$) Faith. ($\uparrow$) Span-Cov. ($\uparrow$) Span-Ext. ($\downarrow$) LEE ($\uparrow$)
Qwen2.5-14B-Instruct
PromptBaseline -0.028 0.74 -0.08 0.60
@@
Table 1: Uncertainty NLE evaluation results across the HealthVer and DRUID datasets (§4.1). For each model (§4.2) we compare PromptBaseline, CLUE-Span, and CLUE-Span+Steering on four metrics: Faith. (§5.1), Span-Cov. (§5.2), Span-Ext. (§5.3), and LEE (§5.4). Bold values mark the best result per metric for each dataset–model pair; “–” indicates inapplicable metrics for PromptBaseline , as it is not supplied with extracted span interactions.
Here, we present the results of our automatic evaluation. For brevity, we refer to Qwen2.5-14B-Instruct, OLMo-2-1124-13B-Instruct, and Gemma-2-9B-it simply as Qwen, OLMo, and Gemma, respectively.
@@
We use Entropy-CCT, a point–biserial correlation bounded by $-1\leq r_{\text{pb}}\leq 1$ (Eq. 9), to measure the faithfulness of NLEs to the model’s uncertainty (§5.1). When $r_{\text{pb}}=0$, the explanation mentions high- and low-impact perturbation words equally often; every $+0.01$ adds roughly one percentage point (pp) to the chance that the explanation names a token that is truly influential for the model’s predictive uncertainty (App. G).
Table 1 shows that PromptBaseline is non-faithful in all six settings with $r_{\text{pb}}$ are all negative values ranging from $-0.03$ to $-0.13$. Thus its NLEs mention truly influential tokens 3–13 pp less often than uninfluential ones—the opposite of faithful behaviour. Both variants of our CLUE reverse this trend. Presenting span interactions in the prompt (CLUE-Span) raises every correlation to non-negative values and peaks at $r_{\text{pb}}=0.089$ on the DRUID–Qwen setting. This means the explanation now mentions about 17 pp more often than PromptBaseline($r_{\text{pb}}=-0.080$). Adding attention steering (CLUE-Span+Steering) lifts the $r_{bp}$ scores to $0.033$ on HealthVer and $0.102$ on DRUID with Qwen model, i.e., net gains of +6 pp and +18 pp over PromptBaseline. Moreover, four of the six positive correlations produced by CLUE-Span+Steering are significant at $p<0.01$ (Table 4 in App. G.3), confirming that the improvements are both substantial and statistically reliable. Particularly large jumps of OLMo on Druid dataset (up to $\Delta r_{\text{pb}}=+0.23\approx+23$ pp) suggest that span-interaction guidance from our CLUE framework is most beneficial for models that initially struggle to align explanations with predictive uncertainty.
@@
We evaluate three futher properties of the generated NLEs: (i) Span-Coverage of extracted conflict-/agreement- span interactions (§5.2), (ii) Span-Extraneous: mention of non-extracted spans (§5.3), and (iii) Label-Explanation Entailment with the generated fact-checking label (§5.4). As Table 1 shows, CLUE-Span+Steering outperforms CLUE-Span in both Span-Coverage and Span-extraneous, consistent with the attention steering method’s effectiveness in directing the model to focus on provided spans during generation (Zhang et al., 2024b). Absolute numbers, however, remain modest (peak Span-Coverage: .44, Span-Extraneous: .20 with Qwen). A Span-Coverage of 1 means the NLE cites every extracted interaction, while a Span-Extraneous score of 0 means it adds none beyond them. This gap highlights considerable headroom for better integrating critical span interactions into the explanations. Among the three backbones, Qwen attains the highest Span-Coverage and the lowest Span-Extraneous scores, a trend that likely reflects its stronger instruction-following ability (see benchmark scores in App. A), and thus larger or more capable models might further narrow the gap. Both variants of our framework achieve stronger label-explanation entailment scores than the baseline, yielding explanations logically consistent with the predicted labels while remaining faithful to the model’s uncertainty patterns (as demonstrated in our faithfulness analysis).
@@ 6.1 Method
We recruited N=12 participants from Prolific (https://www.prolific.com/) to rank explanations generated by PromptBaseline, CLUE-Span, CLUE-Span+Steering for 40 instances (20 from DRUID, 20 from HealthVer) (see details about participants and set-up in App. H.1). Adapting Atanasova et al. (2020), participants ranked explanations in descending order (1st, 2nd, 3rd) according to five criteria, complementary to our automatic evaluation metrics:
  • Helpfulness. The explanation offers information that aids readers to judge the claim and fact-check.
@@
  • Consistency. The explanation contains logically consistent statements to the input, disticnt from Label-Explanation Entailment (§5.4), which measures label-explanation alignment.
  • Overall Quality. Ranking of explanations by their overall quality, considering all criteria above.
@@ 6.2 Results
The results of our evaluation results are depicted in Table 2. Annotator agreement was moderate to low (see App. H.2.1), which we attribute to the relative complexity of the task and individual differences in how the information was perceived.
PromptBase CLUE-S CLUE-SS
@@
Overall 2.025 1.892 1.867
DRUID 1.9 1.917 1.767
HealthVer 2.15 1.867 1.967
@@
Overall 1.875 1.783 1.817
DRUID 1.717 1.75 1.617
HealthVer 2.033 1.817 2.017
@@
Overall 2.05 1.908 1.833
DRUID 1.983 1.983 1.683
HealthVer 2.117 1.833 1.983
@@
Overall 1.967 1.775 1.758
DRUID 1.767 1.75 1.617
HealthVer 2.167 1.8 1.9
@@
Overall 1.967 1.908 1.925
DRUID 1.9 1.9 1.817
HealthVer 2.033 1.917 2.033
Table 2: Mean Average Rank (MAR) for the five human-evaluation criteria applied to explanations from Qwen2.5-14B-Instruct on the HealthVer and DRUID datasets (chosen for its high faithfulness; see §5.5). PromptBaseline, CLUE-Span (CLUE-S), and CLUE-Span+Steering (CLUE-SS) are compared. Lower MAR means a better (higher) average rank; the best score in each row is boldfaced.
The explanations generated by CLUE were preferred by our participants to those generated using PromptBaseline: the explanations generated by CLUE-Span+Steering were rated as most helpful, highest coverage, and containing the least amount of redundant information, while those from CLUE-Span were judged to have the highest consistency and overall quality. Although CLUE-Span+Steering achieves the highest faithfulness (see §5.5), our participants judged its overall quality slightly lower than that of CLUE-Span. A possible reason for this is that although CLUE-Span+Steering adheres closely to the top-$K{=}3$ extracted span interactions (as reflected in its higher Span-Coverage and lower Span-Extraneous scores), it may produce explanations that are slightly less internally consistent or fluent. In contrast, CLUE-Span is less faithful to those extracted spans, but may capture additional points that study participants deemed important, likely because the spans identified as important for model do not fully overlap with those identified by humans Ray Choudhury et al. (2023), highlighting the well-documented trade-off between faithfulness and plausibility Agarwal et al. (2024). Future work on improving the plausibility of the span interactions while retaining their faithfulness may therefore improve the human evaluation scores for CLUE-Span+Steering.
@@ 7 Conclusion
We present the first framework, CLUE, for generating NLEs of model uncertainty by referring to the conflicts and agreements between claims and multiple pieces of evidence in a fact-checking task. Our method, evaluated across three language models and two datasets, demonstrates significant improvements in both faithfulness to model uncertainty and label consistency compared to standard prompting. Evaluations by human participants further demonstrate that the explanations generated by CLUE are more helpful, more informative, less redundant, and more logically consistent with the input. This work establishes a foundation for explainable fact-checking systems, providing end users (e.g., fact-checkers) with grounded, faithful explanations that reflect the model’s uncertainty.
@@
Our paper proposes a novel framework for generating NLEs towards the model’s uncertainty by explicitly pointing to the conflicts or agreements within the claim and multi-evidence interactions. While our framework demonstrates improved explanation quality through rigorous evaluation across three language models and two datasets, we acknowledge several limitations that present opportunities for future research.
Our experiments are constrained to medium-sized models (Qwen2.5-14B-Instruct, Gemma2-9B-it, and OLMo2-13B-Instruct) which were selected based on computational limitations. Although these models show significant improvements over baseline performance, our results suggest that larger models (e.g., 70B parameter scale) with enhanced instruction-following and reasoning capabilities might further improve explanation quality — particularly for coverage and redundancy metrics. Our framework’s modular design readily accommodates such scaling.
In this study we focus on the HealthVer and DRUID datasets, in which claims are paired with discrete pieces of evidence, ideal for studying evidence-conflict scenarios. Future work could investigate more complex evidence structures (e.g., long-form documents), diverse fact-checking sources, and scenarios with more than two pieces of evidence per claim to better reflect real-world fact-checking challenges.
While our evaluation with laypeople confirms that our framework produces explanations of higher quality than prompting, expert evaluations (e.g., with professional fact-checkers) are needed to assess practical utility in high-stakes settings.
@@
This work concerns automated fact-checking, which aims to reduce the harm and spread of misinformation, but nevertheless has the potential for harm or misuse through model inaccuracy, hallucination, or deployment for censorship. Our current work aims to provide explanation that allow users to examine the outputs of these systems more critically, and so we do not see any immediate risks associated with it.
Our work is limited to examining claims, evidence, and explanations in English, and so our results may not be generalisable to other languages. As the task involved complex reasoning about technical subjects, we screened our participants to be native English speakers to ensure that they could fully understand the material and increase the chances of high-quality responses (see H.1 for details). However, this criteria may also introduce or reinforce existing biases and limit the generalisability of our findings. Participants were informed about the study and its aims before agreeing to provide informed consent. No personal data was collected from participants and they received fair payment for their work (approximately 9 GBP/hour).

### v2

@@ 4.1 Faithfulness
To assess whether the NLEs produced by CLUE are faithful to the model’s uncertainty, we adapt the Correlational Counterfactual Test (CCT) Siegel et al. (2024) and propose Entropy-CCT. Entropy-CCT retains CCT’s perturb-and-correlate design, but replaces model’s answer probability-shift (e.g., TVD) with predictive entropy change to directly target uncertainty, which is the object our NLEs aim to explain (see the App. F.2 for comparison).
Following Siegel et al. (2024), for each instance $X$ we construct $n$ perturbations $X^{\prime}$ by inserting a random modifier token (adjective before a noun, or adverb before a verb; App. F.1). Let $u(X)$ denote the model’s uncertainty score defined by Eq. 2. We measure the perturbation’s impact on uncertainty using Absolute Entropy Change (AEC):
$\Delta u(X)=\left|u(X)-u(X^{\prime})\right|.$ (8)
For each perturbation $X^{\prime}$, we also record whether the inserted word appears in the generated NLE, using its presence as a proxy for importance. This yields a binary mention flag $m\in\{0,1\}$ (Siegel et al., 2024; Atanasova et al., 2023).
Let $D_{m}$ denote the set of perturbed examples where the NLE mentions the inserted word and $D_{\lnot m}$ the complementary set where it does not. We compute the point-biserial correlation $r_{\text{pb}}$ (Tate, 1954) between the continuous variable $\Delta u$ and the binary flag $m$ over all perturbations. The Entropy-CCT statistic is:
$\scriptstyle\text{CCT}_{\text{entropy}}=r_{\text{pb}}=\frac{\mathbb{E}_{m}[\Delta u]-\mathbb{E}_{\lnot m}[\Delta u]}{\mathrm{Std}(\Delta u)}\cdot\sqrt{\frac{|D_{m}|\cdot|D_{\lnot m}|}{(|D_{m}|+|D_{\lnot m}|)^{2}}}.$ (9)
Here, $\mathbb{E}_{m}[\Delta u]$ and $\mathbb{E}_{\lnot m}[\Delta u]$ are the mean AEC values for the two groups, and $\mathrm{Std}(\Delta u)$ is the standard deviation across all perturbations. Higher Entropy-CCT indicates that the NLE more often mentions tokens whose insertion induces larger changes in predictive uncertainty, and thus is more faithful to reflect the model’s uncertainty.
@@ 4.2 Span-Coverage
$\text{Span-Coverage}\;=\;\frac{\lvert S_{\text{NLE}}\cap S_{R}(k)\rvert}{\lvert S_{R}(k)\rvert}$. (10)
@@ 4.3 Span-Extraneous
$\text{Span-Extraneous}\;=\;\frac{\lvert S_{\text{NLE}}\setminus S_{R}(k)\rvert}{\lvert S_{\text{NLE}}\rvert}$. (11)
@@ 4.4 Label-Explanation Entailment
We evaluate the extent to which the uncertainty explanation agrees with the model’s predicted label by formulating the task as a natural-language inference (NLI) problem. First, we convert the predicted label into a hypothesis using the template “The claim is supported by / refuted by / neutral to the evidence.” The generated explanation serves as the premise. The resulting premise–hypothesis pair is fed to a widely used off-the-shelf language-inference model, DeBERTa-v3 (He et al., 2023). The Label-Explanation Entailment (LEE) score is the proportion of examples for which the NLI model predicts entailment.
@@ 4.5 Results
Model Method HealthVer DRUID
Faith $\uparrow$ Cov. $\uparrow$ Ext. $\downarrow$ LEE. $\uparrow$ Faith. $\uparrow$ Cov. $\uparrow$ Ext. $\downarrow$ LEE. $\uparrow$
Qwen2.5-14B PromptBaseline -0.028 n/a n/a 0.74 -0.080 n/a n/a 0.60
@@
Table 1: Automatic uncertainty-NLE evaluation. Higher is better for Faith.(§4.1)/Cov. (§4.2)/LEE. (§4.4) and lower is better for Ext. (§4.3). n/a indicates Cov./Ext. are undefined for PromptBaseline because no reference spans are provided. Best results per metric for each dataset-model pair are in bold.
Here, we present the results of our automatic NLE evaluation for the scenario where the input consists of one Claim and two pieces of Evidence [^3] See experiments with three pieces of evidence in App. H.1. For brevity, we refer to Qwen2.5-14B-Instruct, OLMo-2-1124-13B-Instruct, and Gemma-2-9B-it simply as Qwen, OLMo, and Gemma, respectively [^4] See experiments on the reasoning model DeepSeek‑R1‑Distill‑Qwen‑14B in App. H.2 .
@@
We use Entropy-CCT, a point–biserial correlation $r_{\text{pb}}$ (Eq. 9) bounded by [1,1][-1,1] to measure the faithfulness of NLEs to the model’s uncertainty (§4.1). When $r_{\text{pb}}=0$, the explanation mentions high- and low-impact perturbation words equally often; every +0.01+0.01 adds roughly one percentage point (pp) to the chance that the explanation names a token that is truly influential for the model’s predictive uncertainty (App. F.3).
Table 1 shows that PromptBaseline is non-faithful in all six settings with all $r_{\text{pb}}$ values negative ranging from 0.03-0.03 to 0.13-0.13. Thus its NLEs mention truly influential tokens 3–13 pp less often than uninfluential ones-the opposite of faithful behaviour. Both variants of our CLUE reverse this trend. Presenting span interactions in the prompt (CLUE-Span) raises every correlation to non-negative values and peaks at $r_{\text{pb}}=0.089$ on the DRUID–Qwen setting. This means the explanation now mentions about 17 pp more often than PromptBaseline ($r_{\text{pb}}=-0.080$). Adding attention steering (CLUE-Span+Steering) lifts the $r_{\text{pb}}$ scores to $0.033$ on HealthVer and $0.102$ on DRUID with Qwen model, i.e., net gains of +6 pp and +18 pp over PromptBaseline. Moreover, four of the six positive correlations produced by CLUE-Span+Steering are significant at $p<0.01$ (Table 7 in App. F.3), confirming that the improvements are both substantial and statistically reliable. Particularly large jumps of OLMo on Druid dataset (up to $\Delta r_{\text{pb}}=+0.23\approx+23$ pp) suggest that span-interaction guidance from our CLUE framework is most beneficial for models that initially struggle to align explanations with its uncertainty.
@@
We evaluate three further properties of the generated NLEs: (i) Span-Coverage of extracted conflict-/agreement- span interactions (§4.2), (ii) Span-Extraneous: mention of non-extracted spans (§4.3), and (iii) Label-Explanation Entailment with the generated fact-checking label (§4.4). As Table 1 shows, CLUE-Span+Steering outperforms CLUE-Span in both Span-Coverage and Span-Extraneous, consistent with the attention steering method’s effectiveness in directing the model to focus on provided spans during generation (Zhang et al., 2024b). Absolute numbers, however, remain modest (peak Span-Coverage: .44, Span-Extraneous: .20 with Qwen). A Span-Coverage of 1 means the NLE cites every extracted interaction, while a Span-Extraneous score of 0 means it adds none beyond them. This gap highlights considerable headroom for better integrating critical span interactions into the explanations. Among the three backbones, Qwen attains the highest Span-Coverage and the lowest Span-Extraneous scores, a trend that likely reflects its stronger instruction-following ability (see benchmark scores in App. A), and thus larger or more capable models might further narrow the gap. Both variants of our framework achieve stronger label-explanation entailment scores than the baseline, yielding explanations logically consistent with the predicted labels while remaining faithful to the model’s uncertainty.
@@ 5.1 Method
We recruited N=12 participants from Prolific (https://www.prolific.com/) to evaluate 120 explanations generated by PromptBaseline, CLUE-Span, CLUE-Span+Steering for 40 unique claims (20 from DRUID, 20 from HealthVer) (see details of participants and setup in App. I.1). Adapting Atanasova et al. (2020), participants ranked explanations in descending order (1st, 2nd, 3rd) according to five criteria, complementary to our automatic evaluation metrics:
  • Helpfulness. The explanation offers information that aids readers to fact-check the claim.
@@
  • Consistency. The explanation contains logically consistent statements to the input, distinct from Label-Explanation Entailment (§4.4), which measures label-explanation alignment.
  • Overall Quality. Ranking of explanations by considering all criteria above.
@@ 5.2 Results
Our evaluation shows that explanations generated by CLUE are consistently preferred to PromptBaseline (Table 2). Similar to prior work that has evaluated NLEs Atanasova et al. (2020); Huang et al. (2024); Solano et al. (2024), annotator agreement was moderate to low (App. I.2.1), which we attribute to the complexity of the task.
PromptBaseline CLUE
@@
Overall 0.281 0.73
DRUID 0.312 0.688
HealthVer 0.25 0.772
@@
Overall 0.290 0.721
DRUID 0.309 0.691
HealthVer 0.27 0.751
@@
Overall 0.252 0.762
DRUID 0.261 0.739
HealthVer 0.242 0.784
@@
Overall 0.275 0.722
DRUID 0.283 0.717
HealthVer 0.266 0.727
@@
Overall 0.325 0.678
DRUID 0.313 0.688
HealthVer 0.336 0.667
Table 2: Proportion of times explanations were ranked as participants’ first preference for each evaluation metric. Explanations generated by Qwen2.5-14B-Instruct (chosen for their high faithfulness; see §4.5)
The explanations generated by our CLUE framework are ranked as participants’ first preference 67-78% of the time, compared to 24-34% of the time for PromptBaseline. The explanations generated by CLUE are rated as most helpful, containing the least amount of redundant information, highest coverage, consistency, and overall quality approximately twice as often as those generated using PromptBaseline.
Although both CLUE-Span and CLUE-Span+Steering outperform PromptBaseline in automatic metrics (§4.5) and are ranked higher than PromptBaseline in the human study (Table 11 in App. I.2.2), we observe a mild faithfulness-plausibility trade-off when comparing the two CLUE variants. CLUE-Span+Steering achieves the highest automatic faithfulness, whereas CLUE-Span is ranked slightly higher in human-perceived overall quality. A possible reason is that CLUE-Span+Steering adheres closely to the top-k=3k{=}3 span interactions (higher Span-Coverage and lower Span-Extraneous), which improves faithfulness but can reduce fluency/naturalness when some extracted spans are fragmentary. In contrast, CLUE-Span may capture additional points that participants deemed important, but not captured by the extracted span interactions (Ray Choudhury et al., 2023). This pattern aligns with the well-documented trade-off between faithfulness and plausibility (Agarwal et al., 2024; Atanasova et al., 2023; Lu and Ma, 2024), and calls for future work to enhance both aspects for improved NLE generation, e.g., by improving the quality of the extracted span interactions.
6 Related Work
6.1 Uncertainty Quantification in LLMs
Recent work estimates LLM uncertainty primarily with white-box, logit-derived measures: predictive entropy of answer distributions (Kadavath et al., 2022; Yang et al., 2025), aggregation across generations (Malinin and Gales, 2021), and response consistency via semantic similarity (Duan et al., 2024; Kuhn et al., 2023; Nikitin et al., 2024). For closed-source models, uncertainty is typically elicited through verbalised confidence (Lin et al., 2022; Mielke et al., 2022), which is overconfident and unreliable (Yona et al., 2024; Tanneru et al., 2024), or approximated by diversity under paraphrased prompts (Zhang et al., 2024a; Chen and Mueller, 2024), which is compute-intensive and conflates prompt noise with model uncertainty. Accordingly, our method focuses on open-source models and adopts predictive entropy, a simple, interpretable, efficient white-box metric computed from answer logits, avoiding prompt-induced noise.
6.2 Linguistic Expressions of Uncertainty
Numerical uncertainty estimates are difficult for end-users to act upon as they do not address sources of uncertainty (Warren et al., 2025). Linguistic expressions of uncertainty, e.g., phrases such as “I’m sure” (Mielke et al., 2022; Tian et al., 2023; Xiong et al., 2024; Ji et al., 2025; Farquhar et al., 2024; Kim et al., 2024), may be more intuitive to understand than numerical ones (Zimmer, 1983; Wallsten et al., 1993; Windschitl and Wells, 1996). However, they are not necessarily faithful reflections of the model’s uncertainty (Yona et al., 2024; Tanneru et al., 2024) risking misleading users (Steyvers et al., 2025). Moreover, they do not explain why the model is uncertain, as our CLUE method does.
6.3 Generating Natural Language Explanations for Fact-Checking
Natural language explanations (NLEs) justify model predictions for lay readers in fact-checking (Wei Jie et al., 2024), complementing summarization and social-media work that provides graphs, manipulated spans as rationales (Ribeiro et al., 2022; Chan et al., 2023; Huang et al., 2025; Sun et al., 2026). In fact-checking, early systems extracted key sentences from supplied articles as explanations (Atanasova et al., 2020); later work improved NLE fluency via post-editing (Jolly et al., 2022) and on data from fact-checking websites (Feher et al., 2025; Raffel et al., 2020; Beltagy et al., 2020). More recent methods use highlight-based explanations as signals to enhance NLE quality, employing techniques such as graph-based modeling (Yuan et al., 2025) and self-refinement (Wang and Atanasova, 2025). Another line of work focuses on few-shot approaches that avoid model supervision, including prompting GPT-3 to produce evidence summaries (Brown et al., 2020; Stammbach and Ash, 2020), inserting a planning step (Zhao et al., 2024), and leveraging retrieval-augmented language models (Zeng and Gao, 2024). However, existing methods are often either unfaithful to model processes (Atanasova et al., 2023; Siegel et al., 2024, 2025) or fail to address model uncertainty (Warren et al., 2025), limiting their practical utility in fact-checking (Schmitt et al., 2024). Recent work in a fact-checking context has shown that explanations of model uncertainty were judged to be more helpful than providing the model verdict and numerical uncertainty alone (Warren et al., 2026), suggesting that generating such explanations may assist people to reason about the reliability of automated fact-checking predictions. Our framework addresses these gaps by explicitly explaining sources of uncertainty in automatic fact-checking, yielding explanations that are both faithful to model uncertainty and provide actionable insights for fact-checkers.
@@ 7 Conclusion
We present the first framework, CLUE, for generating NLEs of model uncertainty by referring to the conflicts and agreements between claims and multiple pieces of evidence in a fact-checking task. Our method, evaluated across three language models and two datasets, demonstrates significant improvements in both faithfulness to model uncertainty and label consistency compared to standard prompting. Evaluations by human participants further demonstrate that the explanations generated by CLUE are more helpful, more informative, less redundant, and more logically consistent with the input. This work establishes a foundation for explainable fact-checking systems, providing end users with grounded, faithful explanations that reflect the model’s uncertainty. Furthermore, our approach may prove useful for future work in a wide range of information-seeking and retrieval-augmented tasks (e.g., question answering) in which explaining uncertainty given conflicting context is critical.
@@
Our paper proposes a novel framework for generating NLEs towards the model’s uncertainty by explicitly pointing to the conflicts or agreements between the claim and multiple pieces of evidence. While our framework demonstrates improved explanation quality through rigorous evaluation across multiple language models and datasets, we acknowledge several limitations.
Our experiments are constrained to medium-sized models (Qwen2.5-14B-Instruct, Gemma2-9B-It, OLMo‑2‑1124‑13B‑Instruct, and DeepSeek‑R1‑Distill‑Qwen‑14B), which are selected based on computational limitations. Although these models show significant improvements over baseline performance, our results suggest that larger models (e.g., 70B parameter scale) may further enhance explanation quality, particularly in terms of coverage and redundancy. Our framework’s modular design readily accommodates such scaling.
CLUE includes an auxiliary relation labeler $L_{\mathrm{rel}}$ that assigns each extracted span interaction one of {agree, disagree, unrelated}. In the main experiments we use GPT-4o for this step to reduce relation-label noise, but this step is optional: $L_{\mathrm{rel}}$ is modular, and a fully open-weight variant (e.g., Qwen2.5-14B-Instruct) yields comparable downstream NLE quality with only a small drop (App. C.2). Reducing this remaining gap while remaining fully open-weight may require improving relation-label prompting (more local context, higher-quality shots) or lightly fine-tuning an open-weight model on a compact relation-judgment set (App. C.4).
Like most multi-stage pipelines, CLUE can occasionally propagate upstream errors: manual inspection reveals failure modes such as fragmentary spans and occasional relation mislabels, which can in turn reduce NLE faithfulness (App. C.3). These issues could be mitigated by encouraging more complete semantic spans during extraction (e.g., mild constraints during community detection), improving relation-label robustness (prompting or lightweight fine-tuning), and adding lightweight generation-time checks that down-weight unreliable span interactions before producing the final explanation (App. C.4).
In this study, we focus on the HealthVer and DRUID datasets, which pair claims with discrete pieces of evidence, making them ideal for studying evidence-conflict scenarios. Our experiments primarily focus on setups where one claim and two or three pieces of evidence are presented. Although we did not explore setups involving more complex claim-evidence interactions, when multiple claims and/or pieces of evidence are introduced, our framework, grounded in pairwise interactions, can seamlessly accommodate more complex scenarios by simply increasing the number of input claim/evidence pieces; no methodological changes are required.
CLUE targets settings where users need to understand why the model is uncertain by grounding uncertainty in concrete claim-evidence and inter-evidence conflicts/agreements. This focus is motivated by qualitative findings that professional fact-checkers want systems to “show the work” behind a decision, especially when evidence-based tensions drive uncertainty, so that users can decide what to verify next and when to trust the output (Warren et al., 2025). While in this paper, we conduct a human study and find that uncertainty explanations generated by our proposed CLUE framework are more helpful than those from the simple prompting method, we do not study whether these uncertainty-oriented explanations are preferable in every context compared to alternative methods for communicating uncertainty (e.g., numeric confidence (Zimmer, 1983; Wallsten et al., 1993; van der Waa et al., 2020; Liu et al., 2020) or verbal hedges (Lin et al., 2022; Mielke et al., 2022; Yona et al., 2024; Kim et al., 2024)). In-situ evaluation with professional fact-checkers, potentially including interface-level ablation studies (e.g., tuning the number of surfaced interactions $k$ and using compact/expandable formats), would be needed to study the effects of uncertainty explanations on efficiency, error detection, and calibrated trust in other contexts.
@@
This work concerns automated fact-checking, which aims to reduce the harm and spread of misinformation, but nevertheless has the potential for harm or misuse through model inaccuracy, hallucination, or deployment for censorship. Our current work aims to provide explanations that allow users to examine the outputs of these systems more critically, and so we do not see any immediate risks associated with it.
Our work is limited to examining claims, evidence, and explanations in English, and so our results may not be generalisable to other languages. As the task involved complex reasoning about technical subjects, we screened our participants to be native English speakers to ensure that they could fully understand the material and increase the chances of high-quality responses (see I.1 for details). However, this criteria may also introduce or reinforce existing biases and limit the generalisability of our findings. Participants were informed about the study and its aims before agreeing to provide informed consent. No personal data was collected from participants and they received fair payment for their work (approximately 9 GBP/hour).